1. Discuss managerial decisions regarding the cost of capital.

2. What is the cost of capital (WACC) for the firm that you are researching? (Student responses can comment on relative risks and costs of capital)
Answer #1

(1): Managerial decisions with regards to cost of capital are decisions related to various aspects of financial management. These decisions pertain to use of financial leverage, capital structure, dividend policy of the company, decisions related to working capital, appraisal of different projects etc. The cost of capital concept is used to make decisions so as to be able to maximize the value of a firm, make optimal capital budgeting decisions, and properly evaluate the financial performance of a company.

(2): The firm that I am researching is GE (General Electric). The company’s WACC is 4.12%.

WACC = E/(E+D) * cost of equity + D/(E+D) * cost of debt * (1-tax rate)

For GE its weight of equity = 74791/(74791+122272) = 0.3795

Weight of debt = 122272/((74791+122272) = 0.6205

Cost of equity will be computed using CAPM model. Cost of equity = risk free rate + beta*(market risk premium) = 1.52% + (1.14*6%) = 8.36%

Cost of debt = interest expense/debt = 2077/122272 = 1.6987%. The company’s tax rate is 10.255%

Thus WACC = 0.3795 * 8.36% + 0.6205*1.6987%*(1-10.255%)

= 4.12%